The volume of a box is 350 cubic inches. What are its dimensions if the length is twice the width and the height exceeds the width by 2 inches. (Show your solution)
Answer by Alan3354(69443)

The volume of a box is 350 cubic inches. What are its dimensions if the length is twice the width and the height exceeds the width by 2 inches.
----------------
w*2w*(w+2) = 350
2w^3 + 4w^2 = 350
w^3 + 2w^2 - 175 = 0
--> 5 by 10 by 7